# From Prompt to Pipeline: Large Language Models for Scientific Workflow Development in Bioinformatics

### v1

@@ Abstract
Purpose: The increasing complexity of bioinformatics data analysis necessitates the use of Scientific Workflow Systems (SWSs) such as Galaxy and Nextflow to support scalable, reproducible, and automated scientific workflows. However, developing and understanding these workflows remains challenging for domain scientists (e.g., Bioinformaticians, Biologists, or Geneticists), particularly those without programming experience. This difficulty arises from the reliance on numerous black-box tools/modules and the complex infrastructure required to work with SWSs. This study investigates whether modern Large Language Models (LLMs) can assist in generating accurate, complete, and usable bioinformatics workflows and identifies the prompting strategies that maximize their effectiveness. The goal is to assess their potential to lower the barrier for domain specialists, enhance reproducibility, and improve the accessibility, usability, and maintainability of bioinformatics workflows.
Methods: We evaluate three state-of-the-art LLMs, GPT-4o, Gemini 2.5 Flash, and DeepSeek-V3, across three research questions: (1) Can LLMs assist in scientific workflow development? (2) How do LLM-generated workflows compare in terms of completeness, correctness, and usability? (3) What prompting strategies best support workflow generation? We select a set of representative workflows across Galaxy (graphical-based) and Nextflow (script-based), covering diverse bioinformatics tasks such as SNP analysis, RNA-seq, DNA methylation, and data retrieval. The generated workflows are manually evaluated by domain experts for correctness, completeness, tool appropriateness, and executability, using established baselines from the Galaxy Training Network (GTN) and nf-core repositories. To ensure consistency and reduce complexity, all models are used with default settings, and prompting strategies, including instruction-only, role-based, and chain-of-thought, are applied uniformly across tasks to maintain comparability while preserving workflow quality.
Results: LLMs demonstrated varying levels of proficiency across tasks. Gemini 2.5 Flash produced the most accurate and user-friendly workflows for Galaxy SWS, while DeepSeek-V3 performed well for Nextflow. GPT-4o performed best with structured prompts that provided explicit context, while DeepSeek-V3 offered rich technical outputs, including command-line instructions and directory structures, albeit with verbosity and occasional inconsistencies. Prompting strategies significantly influenced output quality, with role-based and chain-of-thought prompts improving completeness and correctness. Overall, LLMs demonstrated promising capabilities in lowering the barrier to scientific workflow development, with prompting strategy and model selection playing critical roles in effectiveness.
Conclusion: This study demonstrates that LLMs can assist scientific workflow development, particularly when guided by well-crafted prompts. By translating natural language into workflow, these models can reduce barriers for both novice and expert users, enhance reproducibility, and accelerate scientific discovery. Future research should expand to other domains, explore adaptive prompting strategies, and develop integrated tooling that leverages LLMs more effectively in real-world scientific environments.
@@
Scientific Workflows, Large Language Models, Bioinformatics, GPT-4o, Gemini 2.5 Flash, DeepSeek-V3, Prompt, Galaxy, Nextflow
@@ 1 Introduction
The exponential growth of biological data and the increasing complexity of computational analyses have made scientific workflows [1] indispensable in modern bioinformatics [2]. These workflows (workflow means scientific workflow for the rest of the paper) are essential for enabling scalable, reproducible, and dependable large-scale data analysis, particularly when deployed on distributed and cloud infrastructures [3, 4]. In bioinformatics, workflows routinely process terabyte-scale datasets generated by advanced DNA, RNA, NGS sequencing, and other technologies [4] across diverse experimental contexts [5]. Typically, these workflows comprise multiple computational steps, including data preprocessing, quality control, sequence aggregation, machine learning for classification or clustering, statistical analysis, and result visualization. Each step is carried out by specific tools/modules, often developed by third parties, facilitated by scientific workflow systems (SWSs)’ no/low-code environments [6]. SWSs have become essential platforms for constructing, executing, and managing bioinformatics workflows by integrating diverse analytical and data processing tools/modules into reproducible pipelines (workflows)[7]. SWSs such as Galaxy [8], Nextflow [9], Snakemake [10], Airflow [11], and others [2, 4, 12, 13] provide structured environments that support the automation, customization, and reuse of complex analyses. In response to the growing demand for scalable and transparent data processing, hundreds of SWSs [14] have emerged, offering access to thousands of bioinformatics tools/modules. For instance, Galaxy’s ToolShed alone hosts over 10,500 tools [15], Nextflow has over 1500 modules, enabling researchers to assemble rich and versatile workflows tailored to their specific needs. These platforms have significantly advanced reproducibility and efficiency in computational biology by facilitating workflow sharing, versioning, and collaboration [16].
@@
Despite the widespread adoption of SWSs, workflow development remains a significant challenge for both novice and expert users [17, 18, 19, 20]. Practitioners, particularly those without programming backgrounds, face steep learning curves due to the need for scripting, tool configuration, dependency management, and infrastructure provisioning. Workflow developers are usually domain scientists with deep subject-matter expertise but limited experience in software engineering or distributed/cloud computing. While SWSs offer a broad array of services, such as data cleaning, statistical analysis, modeling, and high-performance computing, these capabilities are typically leveraged effectively only by a minority of users who are proficient in workflow composition, programming, and infrastructure management [21]. Consequently, many researchers struggle to design workflows to manage complex data and operate on these SWSs. These difficulties stem from the diversity of tools/modules, scripts, and data formats essential to ensure data integrity, reproducibility, and scalability. The resulting complexity hampers workflow development, adaptation, and interpretation, ultimately impeding data exploration and scientific discovery [20, 22].
@@
Recent advancements in LLMs such as GPT-4o [23], Gemini 2.5 Flash [24], and DeepSeek-v3 [25] have demonstrated their exceptional ability to comprehend complex instructions and generate accurate, context-aware code, making them increasingly relevant in software engineering, particularly in code generation [26]. With their enhanced language comprehension and generative capabilities, they are being extensively explored to automate various software design and development aspects [27]. Despite these promising developments, research exploring the application of LLMs in workflows, especially through practical, real-world scenarios in bioinformatics, remains sparse. This gap highlights the need for studies investigating how LLMs can support practitioners in workflow development. Given their ability to generate code, we posit that LLMs can similarly generate accurate workflows with effective prompts. Thus, we plan to investigate whether LLMs can assist in workflow development. We choose GPT-4o, Gemini 2.5 Flash, and DeepSeek-v3 as they are among the most advanced, publicly accessible LLMs as of 2025. GPT-4o is known for its strong general-purpose capabilities, broad tool knowledge, and superior performance in code generation tasks. Gemini 2.5 Flash has demonstrated strengths in structured reasoning, contextual understanding, and system integration, making it a promising candidate for guiding workflow development. DeepSeek-v3, a newer open-source model trained with a code-heavy corpus, shows competitive performance in scientific and software engineering tasks, and offers an opportunity to evaluate performance beyond proprietary models.
@@
Integrating LLMs into bioinformatics workflow development presents a compelling opportunity to reduce the cognitive and technical burden on workflow developers. With appropriate prompting, LLMs could potentially help users generate workflows, configure parameters, select tools/modules, and even interpret error messages. However, concerns persist regarding the reliability of generated outputs, especially in scientific contexts where precision and reproducibility are critical. To ensure the effectiveness of our approach, we have meticulously applied state-of-the-art prompt engineering strategies, drawing on insights from prominent studies in the field [28, 29, 30, 31, 32, 33, 34].
In this study, we conduct a comprehensive evaluation of state-of-the-art LLMs (GPT-4o, Gemini 2.5 Flash, and DeepSeek-V3) in their ability to support scientific workflow development in bioinformatics. We begin by assessing the capacity of these models to explain foundational concepts related to scientific workflows and SWSs, followed by an in-depth evaluation of their ability to develop workflows across two widely used platforms: Galaxy and Nextflow. We then analyze the generated workflows for correctness, completeness, and usability, comparing them against community-curated baselines (Galaxy Training Network [35] and nf-core [36]). Furthermore, we investigate how different prompting strategies (e.g., instruction-only, role-based, and chain-of-thought) affect output quality and workflow fidelity. By identifying recurring challenges and model-specific limitations, we propose directions for improving LLM-supported workflow development, including integrating domain-specific knowledge and refining prompt engineering techniques. This work contributes to the emerging intersection of natural language processing and computational bioinformatics by offering practical insights into how LLMs can bridge the gap between domain scientists and automated workflow generation. Specifically, we aim to answer the following three research questions (RQs):
@@ 2.1 Scientific Workflows
A scientific workflow is a structured and reproducible sequence of data processing tasks, typically modeled as a directed acyclic graph (DAG), where nodes represent computational operations and edges indicate data dependencies [1, 37]. It automates complex scientific procedures, such as data acquisition, transformation, and analysis, thereby accelerating discovery through efficient and reliable execution [38]. Analogous to a cooking recipe, where ingredients, preparation steps, and instructions combine to produce a consistent outcome, scientific workflows ensure reproducibility and scalability in computational experiments. Lin et al. [39] defined workflows as the computerized automation of scientific processes that streamline both tasks and dataflows. In this paper, we use the terms workflow, scientific workflow, and pipeline interchangeably to refer to such structured processes. A workflow can be graphic-based and script-based. Figure 1 illustrates an example of an RNA sequencing workflow steps developed using Nextflow SWS, obtained from [36]. This pipeline takes FASTQ files as input, performs quality control and trimming, aligns or pseudo-aligns reads to the reference genome, quantifies gene and transcript expression, and generates a comprehensive expression matrix alongside detailed QC reports.
@@ 2.3 Large Language Models
LLMs are advanced AI systems capable of processing and generating human-like text using vast training data. Their development has been revolutionized by the transformer architecture introduced by Vaswani et al. in Attention Is All You Need [42]. Leveraging neural networks and transformer variants, LLMs like GPT-3, GPT-4, Codex (OpenAI), DeepSeek-V3 (DeepSeek), Gemini, PaLM (Google), LLaMA (Meta), and Claude (Anthropic) and others [43] have become cornerstones of natural language processing. Tools like Codex and ChatGPT assist developers by suggesting code snippets, completing functions, and enhancing productivity [44]. They aid in code review, bug detection, and quality assurance, promoting reliable software [45]. Additionally, they generate and update documentation, user guides, and API references, simplifying software maintenance [46, 47]. LLMs even support cross-platform development by translating code between languages, improving collaboration among diverse teams [48, 49]. LLMs are widely acknowledged for their ability to generate commonsense knowledge, leveraging the vast training data [3]. This capability makes them practical and accessible knowledge repositories that can be queried through natural language prompts. A key limitation of LLMs is their tendency to produce hallucinations [50, 51].
@@ 2.4 Usages of LLMs
Recent advances in LLMs have sparked significant interest in automating various aspects of software development, including code generation, refactoring, and debugging. Numerous studies have evaluated LLMs for their ability to generate syntactically correct and semantically relevant code from natural language prompts [52, 53, 54]. Beyond code generation, LLMs have also shown promise in interactive programming tasks such as modifying code based on user intent, understanding incomplete problem statements, and suggesting appropriate development patterns [55]. LLMs like Codex, ChatGPT, and GitHub Copilot have demonstrated practical utility in supporting programmers by reducing the need for frequent online searches and offering useful scaffolds for common tasks. However, users often encounter challenges related to debugging, editing, and verifying model-generated code, particularly for more complex programming tasks [56]. Despite these limitations, studies show LLMs can be beneficial in educational settings, improving student learning outcomes and reducing cognitive load during programming exercises [57]. In comparative evaluations, ChatGPT has outperformed other code generation tools in terms of correctness and technical debt, highlighting the growing maturity of general-purpose LLMs in software development [58].
While LLMs have been widely studied in the context of general software engineering, their application in scientific data analysis and workflow automation, particularly in domains like bioinformatics, remains relatively nascent. Initial efforts by Liu et al. [59] explored how code generators like Codex could assist non-programmers in performing data analysis tasks. Similarly, Sänger et al. [3] conducted a qualitative assessment of ChatGPT’s effectiveness in supporting scientific workflow design. More recently, Xu et al. [6] introduced LLM4Workflow, a framework that integrates LLMs into a workflow model. Building on these foundational studies, our research investigates the underexplored potential of LLMs in the domain of bioinformatics workflow development. Specifically, we assess how state-of-the-art LLMs can generate, explain, and refine workflows for scientific tasks using two widely adopted SWSs: Galaxy and Nextflow. By systematically analyzing the influence of different prompting strategies, such as instruction-only, role-based, and chain-of-thought, we aim to unlock more effective and reliable uses of LLMs in enabling non-programmers and domain scientists to construct robust, reproducible, and scalable workflows. This work contributes to bridging the gap between advanced language models and practical scientific computing applications.
@@ 2.5 Prompt Engineering
Prompt engineering refers to the process of crafting input instructions that effectively guide an LLM to generate desired outputs. A prompt typically consists of natural language text that conveys a task, includes relevant context, and may specify output format or style [33]. Since LLMs are sensitive to how information is framed, the structure, clarity, and specificity of a prompt significantly influence the relevance, accuracy, and usefulness of the model’s response [60]. Prompt engineering has therefore emerged as a critical practice for aligning LLM behavior with user goals, especially in complex domains such as scientific workflow development.
In the context of this study, prompt engineering plays a pivotal role in enabling LLMs to understand domain-specific requirements in bioinformatics and translate them into structured workflows. Drawing from recent research that emphasizes prompt design as a programming paradigm in itself [61], we explore strategies [34, 62, 33, 29, 28, 30, 31, 32, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73] for adapting prompts to scientific tasks, such as generating analysis pipelines or interpreting bioinformatics procedures. These strategies include varying prompt templates, combining task descriptions with expected outcomes, and incorporating domain-relevant terminology to improve performance across different use cases [74].

### v2

@@ Abstract
Scientific Workflow Systems (SWSs) such as Galaxy and Nextflow are essential for scalable, reproducible, and automated bioinformatics analyses. However, developing and understanding scientific workflows remains challenging for many domain scientists due to the complexity of tool/module selection, infrastructure requirements, and limited programming expertise. This study explores whether state-of-the-art Large Language Models (LLMs) such as GPT-4o, Gemini 2.5 Flash, and DeepSeek-V3 can assist in generating accurate, complete, and usable bioinformatics workflows. We evaluate a set of representative workflows covering tasks such as RNA-seq, SNP analysis, and DNA methylation across both Galaxy (graphical) and Nextflow (script-based) platforms. To simulate realistic usage, we adopt a tiered prompting strategy: each workflow is first generated using an instruction-only prompt; if the output is incomplete or incorrect, we escalate to a role-based prompt, and finally to chain-of-thought prompting if needed. The generated workflows are evaluated against community-curated baselines from the Galaxy Training Network (GTN) and nf-core, using criteria including correctness, completeness, tool appropriateness, and executability. Results show that LLMs exhibit strong potential in workflow development. Gemini 2.5 Flash produced the most accurate and user-friendly workflows in Galaxy, while DeepSeek-V3 excelled in Nextflow pipeline generation. GPT-4o performed nicely with structured prompts. Prompting strategy significantly influenced output quality, with role-based and chain-of-thought prompts enhancing correctness and completeness. Overall, LLMs can reduce the cognitive and technical barriers to workflow development, making SWSs more accessible to novice and expert users. This work highlights the practical utility of LLMs and provides actionable insights for integrating them into real-world bioinformatics workflow design.
@@
Scientific Workflows , Scientific Workflow Systems , Large Language Models , Bioinformatics , Prompting Techniques , Natural Language to Workflow
journal: Journal of Systems and Software
[…]
organization=Department of Computer Science, University of Saskatchewan, addressline=110 Science Pl, city=Saskatoon, postcode=S7N 5A2, state=SK, country=Canada
@@ 1 Introduction
The exponential growth of biological data and the increasing complexity of computational analyses have made scientific workflows [1] indispensable in modern bioinformatics [2]. These workflows (workflow means scientific workflow for the rest of the paper) are essential for enabling scalable, reproducible, and dependable large-scale data analysis, particularly when deployed on distributed and cloud infrastructures [3, 4]. In bioinformatics, workflows routinely process terabyte-scale datasets generated by advanced DNA, RNA, NGS sequencing, and other technologies [4] across diverse experimental contexts [5]. Typically, these workflows comprise multiple computational steps, including data preprocessing, quality control, sequence aggregation, machine learning for classification or clustering, statistical analysis, and result visualization. Each step is carried out by specific tools/modules, often developed by third parties, facilitated by scientific workflow systems (SWSs)’ no/low-code environments [6]. SWSs have become essential platforms for constructing, executing, and managing bioinformatics workflows by integrating diverse analytical and data processing tools/modules into reproducible pipelines (workflows) [7]. SWSs such as Galaxy [8], Nextflow [9], Snakemake [10], Airflow [11], and others [2, 4, 12, 13] provide structured environments that support the automation, customization, and reuse of complex analyses. In response to the growing demand for scalable and transparent data processing, hundreds of SWSs [14] have emerged, offering access to thousands of bioinformatics tools/modules. For instance, Galaxy’s ToolShed  [^1] https://toolshed.g2.bx.psu.edu/ alone hosts over 10,600 tools, Nextflow has over 1550 modules  [^2] https://nf-co.re/modules/, enabling researchers to assemble rich and versatile workflows tailored to their specific needs. These platforms have significantly advanced reproducibility and efficiency in computational biology by facilitating workflow sharing, versioning, and collaboration [15].
@@
Despite the widespread adoption of SWSs, workflow development remains a significant challenge for both novice and expert users [16, 17, 18, 19]. Practitioners, particularly those without programming backgrounds, face steep learning curves due to the need for scripting, tool configuration, dependency management, and infrastructure provisioning [20]. Workflow developers are usually domain scientists with deep subject-matter expertise but limited experience in software engineering or distributed/cloud computing. While SWSs offer a broad array of services, such as data cleaning, statistical analysis, modeling, and high-performance computing, these capabilities are typically leveraged effectively only by a minority of users who are proficient in workflow composition, programming, and infrastructure management [21]. Consequently, many researchers struggle to design workflows to manage complex data and operate on these SWSs. These difficulties stem from the diversity of tools/modules, scripts, and data formats essential to ensure data integrity, reproducibility, and scalability. The resulting complexity hampers workflow development, adaptation, and interpretation, ultimately impeding data exploration and scientific discovery [19, 22].
@@
Recent advancements in LLMs such as GPT-4o [23], Gemini 2.5 Flash [24], and DeepSeek-v3 [25] have shown strong capabilities in understanding complex instructions and generating accurate, context-aware code, making them highly relevant for code generation and software development tasks [26, 27]. However, their application in real-world bioinformatics workflow development remains underexplored. This gap underscores the need to assess how LLMs can support practitioners in developing scientific workflows. Given their code generation abilities, we posit that LLMs can produce accurate workflows with appropriate prompting. We investigate this potential using three state-of-the-art LLMs. GPT-4o offers broad tool knowledge and excels in general-purpose coding tasks; Gemini 2.5 Flash demonstrates strengths in structured reasoning and system integration; and DeepSeek-v3, a code-focused open-source model, provides a valuable perspective beyond proprietary alternatives.
@@
Integrating LLMs into bioinformatics workflow development presents a compelling opportunity to reduce the cognitive and technical burden on workflow developers. With appropriate prompting, LLMs could potentially help users generate workflows, configure parameters, select tools/modules, and even interpret error messages. However, concerns persist regarding the reliability of generated outputs, especially in scientific contexts where precision and reproducibility are critical. To ensure the effectiveness of our approach, we have meticulously applied state-of-the-art prompt engineering strategies, drawing on insights from prominent studies in this field [28, 29, 30, 31, 32, 33, 34].
In this study, we conduct a comprehensive evaluation of state-of-the-art LLMs (GPT-4o, Gemini 2.5 Flash, and DeepSeek-V3) in their ability to support workflow development in bioinformatics. We begin by assessing the capacity of these models to explain foundational concepts related to scientific workflows and SWSs, followed by an in-depth evaluation of their ability to develop workflows across two widely used platforms: Galaxy and Nextflow. We then analyze the generated workflows for correctness, completeness, and usability, comparing them against community-curated baselines (Galaxy Training Network [35] and nf-core [36]). Furthermore, we investigate how different prompting strategies (e.g., instruction-only, role-based, and chain-of-thought) affect output quality. By identifying recurring challenges and model-specific limitations, we propose directions for improving LLM-supported workflow development, including integrating domain-specific knowledge and refining prompt engineering techniques. This work contributes to the emerging intersection of natural language processing and computational bioinformatics by offering practical insights into how LLMs can bridge the gap between domain scientists and automated workflow generation. Specifically, we aim to answer the following three research questions (RQs):
  • 1.
@@ 2.1 Scientific Workflows
A scientific workflow is a structured and reproducible sequence of data processing tasks, typically modeled as a directed acyclic graph (DAG), where nodes represent computational operations and edges indicate data dependencies [1, 37]. It automates complex scientific procedures, such as data acquisition, transformation, and analysis, thereby accelerating discovery through efficient and reliable execution [38]. Analogous to a cooking recipe, where ingredients, preparation steps, and instructions combine to produce a consistent outcome, scientific workflows ensure reproducibility and scalability in computational experiments. In this paper, we use the terms workflow, scientific workflow, and pipeline interchangeably to refer to scientific workflow. A workflow can be graphic-based and script-based. Figure 1 illustrates an example of an RNA sequencing workflow steps developed using Nextflow SWS, obtained from [36]. This pipeline takes FASTQ files as input, performs quality control and trimming, aligns or pseudo-aligns reads to the reference genome, quantifies gene and transcript expression, and generates a comprehensive expression matrix alongside detailed QC reports.
@@ 2.3 Large Language Models
LLMs are advanced AI systems capable of processing and generating human-like text using vast training data. Their development has been revolutionized by the transformer architecture introduced by Vaswani et al. in Attention Is All You Need [42]. Leveraging neural networks and transformer variants, LLMs like GPT-3, GPT-4, Codex (OpenAI), DeepSeek-V3 (DeepSeek), Gemini, PaLM (Google), LLaMA (Meta), and Claude (Anthropic) and others [43] have become cornerstones of natural language processing. Tools like Codex and ChatGPT assist developers by suggesting code snippets, completing functions, and enhancing productivity [44]. They aid in code review, bug detection, and quality assurance, promoting reliable software [45]. Additionally, they generate and update documentation, user guides, and API references, simplifying software maintenance [46, 47]. LLMs even support cross-platform development by translating code between languages, improving collaboration among diverse teams [48, 49]. LLMs are widely acknowledged for their ability to generate commonsense knowledge, leveraging the vast training data [3]. This capability makes them practical and accessible knowledge repositories that can be queried through natural language prompts. LLMs are also increasingly used in bioinformatics for tasks like protein structure prediction, genomic sequence analysis, drug discovery, gene expression profiling, and pathway analysis, leveraging their ability to model complex biological data with transformer-based architectures [50]. A key limitation of LLMs is their tendency to produce hallucinations [51, 52].
@@ 2.4 Usages of LLMs
Recent advances in LLMs have driven interest in automating software development tasks such as code generation, refactoring, and debugging. Studies show that models like Codex, ChatGPT, and GitHub Copilot can generate syntactically correct and semantically meaningful code from natural language prompts [53, 54, 55], and also assist with interactive tasks like modifying code based on user intent or incomplete specifications [56]. They help reduce reliance on online searches and offer scaffolds for common tasks, though challenges persist in debugging and verifying complex outputs [57]. Despite these issues, LLMs have proven effective in many software development contexts [58, 59].
While LLMs have been widely studied in the context of general software engineering, their application in scientific data analysis and workflow automation, particularly in domains like bioinformatics, remains relatively nascent. Initial efforts by [60] explored how code generators like Codex could assist non-programmers in performing data analysis tasks. Similarly, [3] conducted a qualitative assessment of ChatGPT’s effectiveness in supporting scientific workflow design. More recently, [6] introduced LLM4Workflow, a framework that integrates LLMs into a workflow model. LLMs like BERT and GPT-series have been increasingly adapted for bioinformatics, leading to the development of specialized models such as DNABERT [61], ProteinBERT [62], DNAGPT [63], and ProtGPT2 [64]. These domain-specific variants are designed to process and interpret the unique linguistic structures of biological sequences, such as DNA, RNA, and proteins by conceptualizing them as natural language. Leveraging transformer-based architectures, these models demonstrate strong performance in a variety of tasks, including gene annotation, protein function prediction, variant calling, and motif discovery. Their ability to handle large-scale, complex, and unstructured biological data makes them powerful tools for uncovering meaningful biological patterns, enhancing data-driven decision-making, and accelerating discoveries in genomics, and proteomics. Building on these foundational studies, our research investigates the underexplored potential of LLMs in the domain of bioinformatics workflow development. Specifically, we assess how state-of-the-art LLMs can generate, explain, and refine workflows for scientific tasks. By systematically analyzing the influence of different prompting strategies, such as instruction-only, role-based, and chain-of-thought, we aim to unlock more effective and reliable uses of LLMs in enabling non-programmers and domain scientists to construct robust, reproducible, and scalable workflows. This work contributes to bridging the gap between advanced language models and practical scientific workflow development.
@@ 2.5 Prompt Engineering
Prompt engineering refers to the process of crafting input instructions that effectively guide an LLM to generate desired outputs. A prompt typically consists of natural language text that conveys a task, includes relevant context, and may specify output format or style [33]. Since LLMs are sensitive to how information is framed, the structure, clarity, and specificity of a prompt significantly influence the relevance, accuracy, and usefulness of the model’s response [65]. Prompt engineering has therefore emerged as a critical practice for aligning LLM behavior with user goals, especially in complex domains such as scientific workflow development. In the context of this study, prompt engineering plays a pivotal role in enabling LLMs to understand domain-specific requirements in bioinformatics and translate them into structured workflows. Drawing from recent research that emphasizes prompt design as a programming paradigm in itself [66], we explore strategies [34, 67, 33, 29, 28, 30, 31, 32, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78] for adapting prompts to scientific tasks, such as generating analysis pipelines or interpreting bioinformatics procedures. These strategies include varying prompt templates, combining task descriptions with expected outcomes, and incorporating domain-relevant terminology to improve performance across different use cases [79].